Use the References to access important values if needed for this question a. How many grams of chromium(I) iodide, CrI3, are present in 4.20 moles of this compound? grams b. How many moles of chromium(III iodide are present in 2.42 grams of t moles Submit Answer Retry Entire Group 4 more group attempts remaining
Answer #1

Ans

(a)mass = moles*molar mass= 4.20* 432.71 g/mol = 1817.382 g

(b) moles = mass/molar mass = 2.42/432.71= 5.592*10^-3 moles